OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (irishboy123):

\[|2x + 6| − 4 = 20\] normal algebra first, move that 4!

OpenStudy (anonymous):

how will I ? @IrishBoy123

OpenStudy (irishboy123):

if it was \(z - 4 = 20\), how would you solve for z??!!

OpenStudy (anonymous):

i dont know

OpenStudy (irishboy123):

if you honestl don't know, you may really struggle with absolute value stuff have another go :p

OpenStudy (anonymous):

i really dont know

OpenStudy (irishboy123):

so if we thought that \[|2x + 6| − 4 + 4= 20+4\] what would you say?!